In a closed economy


Y= c+i+g


C= 900+0.5(y-t)


I= 700-30r


T= 800


G= 1200


Md=Ms


Ms=1500


Mt=0.7y


Msp= -80r



Derive the IS and LM equations of the economy.( Express y as a function of r and assume p is fixed at 1.0)



Calculate the short run equilibrium values of y and r in the economy.

1
Expert's answer
2022-03-15T10:29:38-0400

IS= 900+0.5(Y-800)+700-30r+1200

Y= 900+0.5Y-400+700-30r+1200

Y=2400+0.5Y-30r

0.5Y=2400-30r

Y=4800-60r

LM=Ms=Md

Md= 0.7y+(-80r)= 0.7Y-0.8r

1500=0.7Y-0.8r

0.7Y= 1500+0.8r

Y= 2142.857+1.143r


At Equilibrium,

LM= IS

2142.857+1.143r=4800-60r

61.134r= 2657.143

r= 43.46%

Y= 2142.857+1.143(43.46)= 2192.53